# Unifying KV Cache Compression for Large Language Models with LeanKV

### v1

@@
Abstract.
Large language models (LLMs) demonstrate exceptional performance but incur high serving costs due to substantial memory demands, with the key-value (KV) cache being a primary bottleneck. Existing KV cache compression methods, including quantization and pruning, struggle with limitations such as uniform treatment of keys and values and static memory allocation across attention heads. To address these challenges, we introduce LeanKV, a unified KV cache compression framework that enhances LLM serving efficiency without compromising accuracy through three innovations: (1) Hetero-KV quantization, which stores keys at a higher precision than values to reflect their greater impact on attention computations; (2) per-head dynamic sparsity, which allocates memory based on token importance per head and per request; and (3) unified KV compression, integrating mixed-precision quantization and selective pruning to enable a smooth tradeoff between model accuracy and memory efficiency. To efficiently support these techniques, LeanKV introduces systems optimizations including unified paging and on-GPU parallel memory management. Implemented on vLLM, LeanKV compresses the KV cache by 3.0×3.0$\times$3.0 × to 5.0×5.0$\times$5.0 × without accuracy loss and up to 11.0×11.0$\times$11.0 × with under 5% accuracy loss, enhancing throughput by 1.9×1.9$\times$1.9 × to 2.5×2.5$\times$2.5 ×, and up to 6.9×6.9$\times$6.9 ×.
*Equal contribution.
1. Introduction
@@
Large language models (LLMs) like GPT (brown2020language, ; achiam2023gpt, ; ouyang2022training, ) and Gemini (team2023gemini, ) have demonstrated significant potential to impact our daily lives, offering promising applications in areas including chatbots (vicuna, ; chiang2024chatbot, ), programming (gitcopilot, ; guo2024deepseek, ), mathematics (shao2024deepseekmath, ; zhong2024evaluation, ) and medical assistance (saab2024capabilities, ; xie2024preliminary, ). Despite their exceptional performance, hosting LLMs is costly due to their large model sizes, demanding extensive hardware resources. Given the pervasive adoption of LLMs, enhancing serving efficiency has become critically important (aminabadi2022deepspeed, ; fang2021turbotransformers, ; kao2023flat, ; shi2023welder, ; zhang2023shepherd, ; han2022microsecond, ; pope2023efficiently, ; li2023alpaserve, ).
@@
LLMs are fundamentally based on the Transformer architecture (vaswani2017attention, ; wolf2020transformers, ), which employs the attention mechanism to generate tokens sequentially, based on the input prompt and previously generated tokens. Within the attention mechanism, each token is transformed into query, key and value feature vectors. The attention scores, derived from the dot product of the query and key vectors, enable the model to weigh the significance of different tokens. Subsequently, the output is computed as the weighted sum of the value vectors, based on the attention scores. To capture diverse relationships between tokens, the attention mechanism is typically divided into multiple heads (vaswani2017attention, ; shazeer2019fast, ; ainslie2023gqa, ). Each head independently computes attention, and their outputs are concatenated and transformed to create a richer representation of the input data
@@
To avoid redundant computation, LLM inference frameworks typically cache intermediate key and value tensors in memory, commonly referred to as the KV cache (kwon2023efficient, ). The size of the KV cache scales linearly with both sequence length and the number of concurrent requests, often comprising over 90% of total memory consumption (kwon2023efficient, ; yu2022orca, ; gao2018low, ; yang2022infless, ). As state-of-the-art models continue to support longer sequences (dubey2024llama, ; jiang2024mixtral, ; liu2024deepseek, ; achiam2023gpt, ; team2023gemini, ; touvron2023llama, ), the KV cache has emerged as a critical bottleneck for LLM serving efficiency. It limits the number of concurrent requests that can fit in memory and increases attention computation latency due to its memory bandwidth-bound nature (dao2022flashattention, ; dao2023flashattention, ).
@@
Recognizing the critical role of the KV cache in limiting serving efficiency, researchers have investigated various compression techniques, primarily focused on quantization (lin2024qserve, ; flashinfer, ; hooper2024kvquant, ) and pruning (zhang2024h2o, ; xiao2023efficient, ; zhang2024pyramidkv, ; ge2023model, ; li2024snapkv, ). Quantization reduces the size of the KV cache by converting floating-point feature values into integers, representing them based on pre-defined intervals. Pruning, on the other hand, reduces the token sequence length by eliminating tokens deemed insignificant based on attention scores. Despite the promise of these techniques, both of them face limitations in efficiency and practicality. First and foremost, existing quantization methods apply the same treatment to key and value vectors of all tokens, disregarding the different roles of keys and values in attention computation, as well as the varying significance of different tokens. Second, current pruning methods statically allocate memory across attention heads, failing to account for the dynamic attention patterns across heads and requests. Moreover, these pruning methods have not been comprehensively evaluated on workloads requiring advanced reasoning, such as programming and mathematics. Lastly, prior work has largely overlooked the potential of unifying quantization and pruning to achieve even greater serving efficiency.
In this paper, we propose LeanKV, a unified KV cache compression framework that addresses the limitations of existing methods and enhances LLM serving efficiency. LeanKV introduces three algorithmic principles :
1. Hetero-KV. LeanKV compresses key and value vectors using heterogeneous precisions, with key vectors stored at higher precision and value vectors at lower precision. This design leverages the observation that key vectors have a greater influence on attention calculations, as they affect the attention scores of all tokens, whereas value vectors only impact the contributions of individual tokens to the attention output.
@@
2. Per-head Dynamic Sparsity. LeanKV dynamically identifies critical tokens on a per-head and per-request basis, adjusting memory allocation accordingly. This design stems from the observation that attention score distributions exhibit dynamic sparsity across both heads and requests. Within a single request, the number of critical tokens varies across heads, and for the same head, the number of critical tokens also varies across different requests.
3. Unified KV Compression. Built on Hetero-KV quantization and per-head dynamic sparsity, LeanKV unifies KV compression strategies by applying mixed-precision quantization and selective pruning on a per-head and per-request basis, with token importance serving as a unified metric. Critical tokens are stored at higher precision, less important tokens are quantized at lower precision, and the least significant tokens are pruned. This integrated approach enables a smooth tradeoff between model accuracy and memory efficiency.
Implementing LeanKV’s algorithmic innovations presents new challenges in system design, particularly regarding memory management. Unlike traditional attention mechanisms that adopt uniform memory allocation across attention heads, LeanKV requires dynamic, per-head memory allocation for each request, significantly increasing system complexity. Moreover, LeanKV must flexibly configure memory to accommodate varying numbers of tokens stored at different precision levels or pruned entirely, further complicating memory management.
@@
To address these challenges, LeanKV introduces two novel techniques. First, LeanKV employs unified paging, which divides memory into fixed-size pages containing vectors stored at various precisions. Each page is dynamically parsed based on the precision configuration required by the head and request, while an attention GPU kernel is co-designed with the page layout to maximize memory bandwidth utilization. Second, LeanKV introduces an on-GPU memory management mechanism that accelerates memory management by parallelizing it directly on the GPU with minimal metadata overhead.
@@
We implement LeanKV on vLLM (kwon2023efficient, ), a state-of-the-art LLM inference system, and evaluate it across three prominent model families using representative workloads, including tasks that demand advanced reasoning such as mathematics and coding. Our experimental results demonstrate that LeanKV offers a superior cost-accuracy tradeoff compared to existing KV cache management methods. Specifically, LeanKV compresses the KV cache by 3.0×3.0$\times$3.0 × to 5.0×5.0$\times$5.0 × without loss in model accuracy, and by up to 11.0×11.0$\times$11.0 × with a marginal accuracy reduction of under 5%, resulting in throughput improvement of 1.9×1.9$\times$1.9 × to 2.5×2.5$\times$2.5 ×, and up to 6.9×6.9$\times$6.9 × correspondingly.
2. Background and Related Work
@@
2.1. Large Language Models
Transformer Architecture. Large language models (LLMs) are predominantly built upon the Transformer architecture (vaswani2017attention, ), which is used to capture complex long-range dependencies in sequential data. At the core of Transformer architecture is the attention mechanism, which allows each token in a sequence to weigh the significance of other tokens when constructing its contextualized representation. Mathematically, the attention computation is defined as:
(1) $\operatorname{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})_{i}$ $=\sum_{j=1}^{i}\operatorname{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}\right)_{ij}\mathbf{v}_{j}$
@@ 2.1. Large Language Models
To capture a broader range of interactions between tokens, Transformer models employ multi-head attention (MHA), where each head independently computes attention using distinct projections of queries, keys, and values. The outputs from different heads are concatenated and linearly transformed to create a more expressive representation. Grouped-query attention (GQA) (ainslie2023gqa, ) improves the efficiency of MHA by allowing a group of query heads to share the same projection of keys and values, referred to as a KV head.
@@
Autoregressive Generation and KV Cache. The autoregressive generation process of LLMs consists of two phases: the prompt phase, during which the model computes latent representations for all tokens in the prompt, the user-provided context, and generates the first new token, and the generation phase, where the model iteratively generates subsequent tokens by attending to both the prompt and previously generated tokens. To avoid redundant computations across generation steps, KV cache is introduced to store the keys and values of all previous tokens. However, the size of the KV cache grows linearly with both the sequence length and batch size, quickly becoming a bottleneck for inference throughput (yu2022orca, ; kwon2023efficient, ). Therefore, efficient KV cache management is critical for alleviating memory bottlenecks and improving LLM serving efficiency.
2.2. KV Cache Optimization
@@
PagedAttention. Static KV cache management systems (yu2022orca, ; FasterTransformer, ) reserve memory for the maximum possible sequence length, leading to considerable memory waste when actual sequences are shorter. To mitigate this inefficiency, vLLM (kwon2023efficient, ) introduces PagedAttention, which partitions the KV cache into pages, each containing a fixed number of tokens, and allocates the pages on-demand as the sequence length grows. By managing the KV cache at the granularity of pages, vLLM reduces memory waste and enables larger batch sizes to improve serving efficiency.
@@
KV Cache Quantization. Quantization (nagel2021white, ; dettmers2022gpt3, ; guo2023olive, ; hooper2024kvquant, ) reduce the KV cache size by approximating high-precision floating points with discrete low-bit values. To quantize a given tensor $\mathbf{X}$, we first calculate the scale $s$ and zero point $z$ as shown in Eq. (2), where $\mathbf{X}_{\text{max}}$ and $\mathbf{X}_{\text{min}}$ represent the maximum and minimum values in $\mathbf{X}$, and $b$ is the bit-width. Then, quantization is applied element-wise as: $\mathbf{Q}=\operatorname{round}\left(\frac{\mathbf{X}+z}{s}\right)$. During inference, the original values are approximately reconstructed via dequantization: $\hat{\mathbf{X}}=s\cdot\mathbf{Q}-z$. Unlike the quantized tensor $\mathbf{Q}$, the metadata $s$ and $z$ are kept in higher precision (e.g., FP16) to ensure more accurate dequantization.
(2) $s=\frac{\mathbf{X}_{\text{max}}-\mathbf{X}_{\text{min}}}{2^{b}-1},\quad z=-\mathbf{X}_{\text{min}}$
@@
State-of-the-art KV cache quantization methods such as Atom (zhao2024atom, ) and Qserve (lin2024qserve, ) apply this process to each key and value vector independently. However, these uniform quantization methods, which apply a fixed bit-width across keys and values of all tokens, may not be optimal. They overlook the varying importance of tokens as well as the different roles of keys and values in the attention calculation.
@@
KV Cache Pruning. Several recent works (zhang2024h2o, ; xiao2023efficient, ; li2024snapkv, ; zhang2024pyramidkv, ; liu2024scissorhands, ; ge2023model, ) have explored KV cache pruning to alleviate the memory bottleneck in LLM inference, which can be considered as an extreme case of quantization. These methods use attention scores to assess token importance and evict less important tokens, thereby reducing the KV cache size. The effectiveness of these methods, however, is limited by their static and inflexible allocation of memory resources. Specifically, H2O (zhang2024h2o, ) and SnapKV (li2024snapkv, ) allocate the same memory budget uniformly across all attention heads and layers. PyramidKV (zhang2024pyramidkv, ) allocates more memory in lower layers and less in higher ones, based on the empirical observation that the number of important tokens shrinks as the model depth increases, but still relies on static heuristics and cannot dynamically adjust cache allocations to suit various workloads. Complementary to KV cache pruning, another line of work (sheng2023flexgen, ; lee2024infinigen, ) addresses GPU memory constraints by offloading the entire KV cache to CPU memory and selectively loading important tokens to the GPU for attention computations. However, these methods do not inherently reduce the KV cache size and incur significant latency due to data transfers between the CPU and GPU.
3. Motivation
In this section, we present the key findings that underpin the design of LeanKV, as well as their implications on KV cache management.
3.1. Impact of Key And Value
@@
(3) $\operatorname{Attn}(\mathbf{Q},\mathbf{K},\mathbf{V})_{i}$ $=\sum_{j=1}^{i}\underbrace{\operatorname{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}\right)_{ij}|\mathbf{v}_{j}|}_{\text{Coefficient}}\underbrace{\frac{\mathbf{v}_{j}}{|\mathbf{v}_{j}|}}_{\text{Unit vector}}$
Figure 1 presents the distributions of the average attention score per token and the norms of value vectors across three representative layers of the Llama3-8B model (dubey2024llama, ), evaluated on the first one thousand articles from the Wikitext dataset (merity2016pointer, ). Notably, the attention scores span seven orders of magnitude, vastly exceeding the range of value vector norms, which only cover at most two orders of magnitude. Similar patterns are also observed in the larger Llama3-70B model. This pronounced disparity highlights the pivotal role of attention scores in determining each token’s contribution to the attention output. We therefore conclude that key vectors exert a broader and more impactful influence than value vectors, motivating further exploration of processing key and value vectors at distinct precision levels.
Figure 1. Distribution of attention score and value vector norm in Llama3-8B.
3.2. Dynamic Sparsity Patterns
Existing approaches to KV cache pruning primarily rely on static memory allocation for each transformer layer in LLMs, either distributing memory uniformly (xiao2023efficient, ; zhang2024h2o, ; ge2023model, ; li2024snapkv, ) or allocating it based on offline profiling (zhang2024pyramidkv, ). However, LLMs exhibit dynamic sparsity patterns across both attention heads and individual requests, which static memory allocation cannot adequately handle. To evaluate the sparsity of attention computation, we analyze the minimum number of critical tokens required to retain the majority of information, specifically by preserving a target percentage (e.g., 95%) of the total attention score.

### v2

@@
Abstract
Large language models (LLMs) exhibit exceptional performance but incur significant serving costs due to their substantial memory requirements, with the key-value (KV) cache being a primary bottleneck. Existing KV cache compression techniques, such as quantization and pruning, apply uniform treatment to both keys and values, and discard unimportant tokens entirely, overlooking the fine-grained differences in significance of various components within the KV cache. To address these limitations, we introduce LeanKV, a framework that advances KV cache compression by exploiting three levels of differentiation in the KV cache: (1) the differing impact of keys and values on attention computation, (2) the varying importance of tokens, and (3) the diverse dynamic sparsity patterns across attention heads. At the core of LeanKV is an on-GPU memory manager that compacts fragmented free memory list into contiguous regions in parallel, effectively translating sparsity in the KV cache into performance gains. We evaluate LeanKV on several mainstream models, including the recent "thinking model". LeanKV is able to compress the KV cache by 2.7×2.7$\times$2.7 × to 5.7×5.7$\times$5.7 × with near-lossless accuracy on complex workloads requiring sophisticated reasoning and long-generation capabilities, and enhances throughput by 1.9×1.9$\times$1.9 × to 5.4×5.4$\times$5.4 ×.
1 Introduction
@@
Large language models (LLMs) like GPT [9, 4, 46] and Gemini [57] have demonstrated significant potential to impact our daily lives, offering promising applications in areas including chatbots [59, 12], programming [21, 24], mathematics [50, 71] and medical assistance [49, 65]. Despite their exceptional performance, hosting LLMs is costly due to their large model sizes, demanding extensive hardware resources. Given the pervasive adoption of LLMs, enhancing serving efficiency has become critically important [6, 18, 31, 53, 68, 25, 47, 36].
@@
LLMs are fundamentally based on the Transformer architecture [61, 62], which employs the attention mechanism to generate tokens sequentially, based on the input prompt and previously generated tokens. Within the attention mechanism, each token is transformed into query, key and value feature vectors. The attention scores, derived from the dot product of the query and key vectors, enable the model to weigh the significance of different tokens. Subsequently, the output is computed as the weighted sum of the value vectors, based on the attention scores. To capture diverse relationships between tokens, the attention mechanism is typically divided into multiple heads [61, 51, 5]. Each head independently computes attention, and their outputs are concatenated and transformed to create a richer representation of the input data.
@@
To avoid redundant computation, LLM inference frameworks typically cache intermediate key and value tensors in memory, commonly referred to as the KV cache [32]. The size of the KV cache scales linearly with both sequence length and the number of concurrent requests, often comprising over 90% of total memory consumption [32, 67, 19, 66]. As state-of-the-art models continue to support longer sequences [17, 30, 38, 4, 57, 60], and with the rise of recent "thinking models" [23, 29, 58] that generate extended reasoning processes, the KV cache has emerged as a critical bottleneck for LLM serving efficiency. It limits the number of concurrent requests that can fit in memory and increases attention computation latency due to its memory bandwidth-bound nature [15, 14].
@@
Recognizing the critical role of the KV cache in limiting serving efficiency, researchers have investigated various compression techniques, primarily focused on quantization [37, 2, 27] and pruning [69, 64, 10, 20, 35]. Quantization reduces the size of the KV cache by converting floating-point feature values into integers, representing them based on pre-defined intervals. Pruning, on the other hand, reduces the token sequence length by eliminating unimportant tokens based on attention scores. While promising, these techniques overlook the fine-grained differentiation in the significance of various KV cache components, limiting their overall effectiveness. First and foremost, existing quantization methods save key and value vectors of all tokens at a uniform precision, disregarding the distinct roles of keys and values and the varying importance of tokens. Second, current pruning methods employ static memory allocation across attention heads, neglecting the dynamic and heterogeneous attention patterns observed across heads and requests. Moreover, these pruning methods have not been comprehensively evaluated on workloads requiring advanced reasoning, such as programming and mathematics.
In this paper, we present LeanKV, a framework that advances KV cache compression by leveraging three levels of differentiation within the KV cache:
1. Differentiated Keys and Values. LeanKV assigns higher precision to key vectors than to value vectors, based on the observation that key vectors play a more significant role in attention computations.
2. Token Importance Differentiation. Guided by attention scores, LeanKV stores tokens at varying precision levels based on their significance. Most critical tokens are quantized at higher precision, less critical tokens are quantized at lower precision, and the least significant tokens are pruned. This hierarchical approach enables a flexible tradeoff between model accuracy and memory efficiency.
@@
3. Per-head Dynamic Sparsity. Attention score distributions exhibit dynamic sparsity across both heads and requests. Within a single request, the number of critical tokens varies across heads, and for the same head, the number of critical tokens also varies across different requests. To address this variability, LeanKV dynamically identifies critical tokens per-head, per-request, adapting memory allocation accordingly.
These differentiations wthin the KV cache pose new challenges in memory management. In LeanKV, each request requires varying amounts of memory for each attention head to store tokens at different precision levels. Moreover, key and value vectors of the same token are compressed at different precisions. These complexities significantly increase memory management overhead, compared to traditional attention mechanisms, where memory allocation is uniform across all heads.
To address these challenges, LeanKV employs an on-GPU memory manager that efficiently handles the irregular memory usage patterns across heads and requests via parallel KV compaction. Rather than performing memory management operations sequentially on CPU for each head and request, parallel KV compaction optimizes memory allocation and recycling by packing fragmented free memory lists into contiguous regions directly on the GPU in parallel. Parallel KV compaction is supported by three key on-GPU data structures:
@@
1. Unified Pages. GPU memory is divided into fixed-size pages containing vectors stored at various precisions. Each page is dynamically parsed according to the precision requirements of the head and request, while an attention GPU kernel is co-designed with the page layout to maximize memory bandwidth utilization.
2. Circular Free Page List. All page IDs are managed in a centralized, GPU-resident circular list. Both free and used pages are kept in contiguous regions, enabling efficient page allocation and recycling via parallel prefix sum.
3. Bidirectional Page Table. Instead of maintaining separate page tables for high- and low-precision pages, LeanKV consolidates both into a single bi-directional page table to minimize metadata overhead. High precision page IDs grow from the left side of the table, and low-precision page IDs grow from the right side.
@@
We implement LeanKV on vLLM [32], a state-of-the-art LLM inference system, and evaluate it across three prominent model families including the emerging thinking model. The evaluation is conducted on representative workloads that demand advanced reasoning capabilities, such as mathematics and coding. Our experimental results demonstrate that LeanKV offers a superior cost-accuracy tradeoff compared to existing KV cache management methods. Specifically, LeanKV compresses the KV cache by 2.7×2.7$\times$2.7 × to 5.7×5.7$\times$5.7 × with near-lossless accuracy, resulting in throughput improvement of 1.9×1.9$\times$1.9 × to 5.4×5.4$\times$5.4 ×. The throughput improvement is primarily enabled by the parallel KV compaction mechanism, which reduces management overhead by three orders of magnitude, decreasing the proportion of memory management time in total execution latency from up to $76\%$ to less than $1\%$ compared to a sequential CPU-based implementation.
2 Background and Related Work
@@
2.1 Large Language Models
Transformer Architecture. Large language models (LLMs) are predominantly built upon the Transformer architecture [61], which is used to capture complex long-range dependencies in sequential data. At the core of Transformer architecture is the attention mechanism, which allows each token in a sequence to weigh the significance of other tokens when constructing its contextualized representation. Mathematically, the attention computation is defined as:
$\operatorname{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})_{i}$ $=\sum_{j=1}^{i}\operatorname{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}\right)_{ij}\mathbf{v}_{j}$ (1)
@@ 2.1 Large Language Models
To capture a broader range of interactions between tokens, Transformer models employ multi-head attention (MHA), where each head independently computes attention using distinct projections of queries, keys, and values. The outputs from different heads are concatenated and linearly transformed to create a more expressive representation. Grouped-query attention (GQA) [5] improves the efficiency of MHA by allowing a group of query heads to share the same projection of keys and values, referred to as a KV head.
@@
Autoregressive Generation and KV Cache. The autoregressive generation process of LLMs consists of two phases: the prompt phase, during which the model computes latent representations for all tokens in the prompt, the user-provided context, and generates the first new token, and the generation phase, where the model iteratively generates subsequent tokens by attending to both the prompt and previously generated tokens. To avoid redundant computations across generation steps, KV cache is introduced to store the keys and values of all previous tokens. However, the size of the KV cache grows linearly with both the sequence length and batch size, quickly becoming a bottleneck for inference throughput [67, 32]. Therefore, efficient KV cache management is critical for alleviating memory bottlenecks and improving LLM serving efficiency.
2.2 KV Cache Optimization
@@
PagedAttention. Static KV cache management systems [67, 1] reserve memory for the maximum possible sequence length, leading to considerable memory waste when actual sequences are shorter. To mitigate this inefficiency, vLLM [32] introduces PagedAttention, which partitions the KV cache into pages, each containing a fixed number of tokens, and allocates the pages on-demand as the sequence length grows. By managing the KV cache at the granularity of pages, vLLM reduces memory waste and enables larger batch sizes to improve serving efficiency.
@@
KV Cache Quantization. Quantization [43, 16, 22, 27] reduce the KV cache size by approximating high-precision floating points with discrete low-bit values. To quantize a given tensor $\mathbf{X}$, we first calculate the scale $s$ and zero point $z$ as shown in Eq. (2), where $\mathbf{X}_{\text{max}}$ and $\mathbf{X}_{\text{min}}$ represent the maximum and minimum values in $\mathbf{X}$, and $b$ is the bit-width. Then, quantization is applied element-wise as: $\mathbf{Q}=\operatorname{round}\left(\frac{\mathbf{X}+z}{s}\right)$. During inference, the original values are approximately reconstructed via dequantization: $\hat{\mathbf{X}}=s\cdot\mathbf{Q}-z$. Unlike the quantized tensor $\mathbf{Q}$, the metadata $s$ and $z$ are kept in higher precision (e.g., FP16) to ensure more accurate dequantization.
$s=\frac{\mathbf{X}_{\text{max}}-\mathbf{X}_{\text{min}}}{2^{b}-1},\quad z=-\mathbf{X}_{\text{min}}$ (2)
@@
State-of-the-art KV cache quantization methods such as Atom [70] and Qserve [37] apply this process to each key and value vector independently. However, these uniform quantization methods, which apply a fixed bit-width across keys and values of all tokens, may not be optimal. They overlook the varying importance of tokens as well as the different roles of keys and values in the attention calculation.
@@
KV Cache Pruning. Several recent works [69, 64, 35, 10, 40, 20] have explored KV cache pruning to alleviate the memory bottleneck in LLM inference, which can be considered as an extreme case of quantization. These methods use attention scores to assess token importance and evict less important tokens, thereby reducing the KV cache size. The effectiveness of these methods, however, is limited by their static and inflexible allocation of memory resources. Specifically, H2O [69] and SnapKV [35] allocate the same memory budget uniformly across all attention heads and layers. PyramidKV [10] allocates more memory in lower layers and less in higher ones, based on the empirical observation that the number of important tokens shrinks as the model depth increases, but still relies on static heuristics and cannot dynamically adjust cache allocations to suit various workloads. Complementary to KV cache pruning, another line of work [52, 33] addresses GPU memory constraints by offloading the entire KV cache to CPU memory and selectively loading important tokens to the GPU for attention computations. However, these methods do not inherently reduce the KV cache size and incur significant latency due to data transfers between the CPU and GPU.
3 The Case for Differentiated KV
We present the case for three levels of differentiation within the KV cache, motivating the design of LeanKV.
3.1 Differentiated Impacts of Keys and Values
@@
$\operatorname{Attn}(\mathbf{Q},\mathbf{K},\mathbf{V})_{i}$ $=\sum_{j=1}^{i}\underbrace{\operatorname{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}\right)_{ij}|\mathbf{v}_{j}|}_{\text{Coefficient}}\underbrace{\frac{\mathbf{v}_{j}}{|\mathbf{v}_{j}|}}_{\text{Unit vector}}$ (3)
Figure 1 presents the distributions of the average attention score per token and the norms of value vectors across three representative layers of the Llama3-8B model [17], evaluated on the first one thousand articles from the Wikitext dataset [42]. Notably, the attention scores span seven orders of magnitude, vastly exceeding the range of value vector norms, which only cover at most two orders of magnitude. Similar patterns are also observed in the larger Llama3-70B model. This pronounced disparity highlights the pivotal role of attention scores in determining each token’s contribution to the attention output. We therefore conclude that key vectors exert a broader and more impactful influence than value vectors, motivating further exploration of processing key and value vectors at distinct precision levels.
Figure 1: Distribution of attention score and value vector norm in Llama3-8B.
3.2 Differentiated Token Importance
Tokens contribute to attention outputs with varying degrees of importance, as reflected by their attention scores. By exploiting these differences, we can apply finer-grained compression strategies that go beyond uniformly quantizing all tokens [70, 37] or exclusively pruning the least important ones [69, 10]. Specifically, we store progressively less important tokens at decreasing precision levels, ultimately leading to the pruning of the least important tokens.
To investigate the feasibility of this hierarchical approach, we experiment with a simple design: quantizing both the key and value vectors of significant tokens to 8 bits while pruning insignificant ones. Figure 2 compares the performance of the FP16 baseline, 8-bit quantization, pruning, and the combination of 8-bit quantization with pruning on the general knowledge question answering benchmark MMLU [26] using the Llama3-8B model. For both pruning and the combined approach, approximately 50% of insignificant tokens are pruned. The resulting accuracy is nearly indistinguishable among the four approaches. However, the combination of 8-bit quantization and pruning achieves the lowest memory usage. The improved efficiency of this combined approach motivates us to further explore the spectrum of KV cache compression by introducing varying levels of precision to accommodate tokens of differing significance.
Figure 2: Accuracy and normalized KV cache memory usage of the Llama3-8B model on MMLU
3.3 Dynamic Sparsity Patterns
Existing approaches to KV cache pruning primarily rely on static memory allocation for each transformer layer in LLMs, either distributing memory uniformly [64, 69, 20, 35] or allocating it based on offline profiling [10]. However, LLMs exhibit dynamic sparsity patterns that vary across both attention heads and requests, which static memory allocation cannot adequately handle. To evaluate the sparsity of attention computation, we analyze the minimum number of critical tokens required to retain the majority of information, specifically by preserving a target percentage (e.g., 95%) of the total attention score.